# Improving Image Classification Accuracy through Complementary Intra-Class and Inter-Class Mixup

### 21 Mar 2024 v1

@@ Abstract
MixUp and its variants, such as Manifold MixUp, have two key limitations in image classification tasks. First, they often neglect mixing within the same class (intra-class mixup), leading to an underutilization of the relationships among samples within the same class. Second, although these methods effectively enhance inter-class separability by mixing between different classes (inter-class mixup), they fall short in improving intra-class cohesion through their mixing operations, limiting their classification performance. To tackle these issues, we propose a novel mixup method and a comprehensive integrated solution. Our mixup approach specifically targets intra-class mixup, an aspect commonly overlooked, to strengthen intra-class cohesion—a feature not provided by current mixup techniques. For each mini-batch, our method utilizes feature representations of unaugmented original images from each class within the mini-batch to generate a single synthesized feature representation through random linear interpolation. All synthesized representations for this mini-batch are then fed into the classification and loss layers to calculate an average classification loss that can markedly enhance intra-class cohesion. Moreover, our integrated solution seamlessly combines our intra-class mixup method with an existing mixup approach, which could be an image-level mixup method like MixUp or a feature-level mixup method like Manifold MixUp. This comprehensive solution incorporates inter- and intra-class mixup in a balanced manner while concurrently improving intra-class cohesion and inter-class separability. Experimental results on six public datasets demonstrate that our integrated solution achieves a 0.1% to 3.43% higher accuracy than the best of either MixUp or our intra-class mixup method, averaging a 1.16% gain. It also outperforms the better performer of either Manifold MixUp or our intra-class mixup method by 0.12% to 5.16%, with an average gain of 1.11%. These findings highlight the effectiveness of our integrated solution in improving classification accuracy. Given that our integrated solution is model-agnostic, it offers significant potential for application in other domains where mixup methods have shown promise, such as speech and text classification. Our source code is publicly available at https://github.com/wxitxy/complementary_intra-class_and_inter-class_mixup.git.
@@ 1 Introduction
Figure 2: t-SNE visualization of feature representations for testing images from the CIFAR-100 dataset: comparing three methods – the baseline (no mixing operation), Manifold MixUp, and our proposed mixup method.
@@
Despite the significant success achieved by existing mixup methods, a closer examination reveals two critical limitations. The first concerns the mixing strategies—namely, intra-class and inter-class mixup. Existing methods such as MixUp, Manifold MixUp, CutMix, and SaliencyMix often neglect intra-class mixup, which involves blending images or hidden representations from the same class. This oversight leads to an underutilization of the relationships among samples within the same class, potentially constraining image classification performance. The root of this issue lies in the shuffle operation; it results in a disproportionately high number of mixed pairs generated from different classes compared to those from the same class. Consequently, there is a predominant emphasis on inter-class mixup, which mixing images or hidden representations from different classes. As illustrated in Figure 1, with 16 classes and a mini-batch size of 32, the average probability of occurrence for intra-class mixup during a shuffle operation is only approximately 9%. If the number of classes and the mini-batch size both increase to 128, this probability decreases to below 2%. In this article, we refer to these techniques that primarily focus on inter-class mixup as “inter-class mixup methods.”
@@
Additionally, the second limitation arises with respect to the mixing outcomes. Existing mixup methods, such as MixUp and Manifold MixUp, fall short in enhancing intra-class cohesion of image feature representations through mixing operations. In this article, the feature representation of an image refers to the input vector that is fed into the neural network’s final classification layer—a fully connected layer responsible for predicting scores for each class. To be specific, the incidental intra-class mixup in current mixup methods indeed has the potential to facilitate the model’s learning of common features among images from the same class, potentially leading to more cohesive clustering of feature representations. Nevertheless, the goal of forming compact class clusters is not explicit, and its effectiveness is significantly diminished due to the rarity of intra-class mixup. On the other hand, the predominant inter-class mixup they implement does enhance the distinction among different classes by promoting smoother predictions between classes but fails to offer any mechanism for reducing intra-class variability. As illustrated in Figure 2, especially within the region enclosed by the dashed blue box, the distinction among classes achieved by Manifold MixUp is noticeably improved compared to the baseline devoid of any mixing operations. However, the feature distribution of each class does not show a marked increase in clustering compared to the baseline. This is particularly obvious in comparison to the enhanced cohesion achieved by our mixup method proposed in this article, as is evident within the red dashed box. In classification tasks, both intra-class cohesion and inter-class separability are known to be vital and equally significant. Consequently, the lack of a dedicated mechanism to improve intra-class cohesion in existing mixup methods may limit their performance in classification tasks.
@@
To overcome these limitations, we propose a novel mixup method that significantly enhances intra-class cohesion through targeted intra-class mixup. Furthermore, we introduce a comprehensive integrated solution that synergistically combines the benefits of both our intra-class mixup method and existing mixup methods, effectively addressing the two critical limitations identified in current mixup practices.
@@
Our proposed mixup method focuses on intra-class mixup, an aspect that existing mixup techniques overlook, with the aim of enhancing intra-class cohesion—a benefit not provided by current mixup techniques. Specifically, during training, our method generates a single synthesized feature representation for each class in every mini-batch. This synthesis is achieved through random linear interpolation among the feature representations of unaugmented original images within the same class. After the synthesized feature representations for all classes in the mini-batch are generated, they are fed into the final classification layer to produce predictions. An intra-class mixup loss is then calculated by averaging the classification losses of these synthesized representations. This loss is designed to promote the model to cluster image feature representations from the same class more tightly, resulting in enhanced intra-class cohesion at the end of training.
@@
Building upon our intra-class mixup method and existing mixup methods, we have developed a comprehensive integrated solution. Our solution leverages an existing mixup method—such as MixUp or Manifold MixUp—as the inter-class mixup component, while our intra-class mixup method serves as the intra-class counterpart. The intra- and inter-class mixup components work together in a complementary manner, with the final loss calculated as a weighted sum of their losses, modulated by a balancing hyperparameter $\beta$. This integrated approach not only uniquely combines both inter- and intra-class mixup strategies in a balanced manner, but also simultaneously improves both intra-class cohesion and inter-class separability. It is capable of adapting to the distinct commonalities within classes and differences among classes that are inherent in various image classification tasks, thereby enhancing the model’s generalization capabilities and robustness. Significantly, our solution allows for flexible substitution of the inter-class mixup component; numerous existing mixup methods, including image-level techniques like MixUp and feature-level techniques like Manifold MixUp, can be implemented as the inter-class mixup component.
@@
It is worth noting that our integrated solution does not deliberately exclude the occasional intra-class mixup that occurs in the existing mixup method as an implementation of the inter-class mixup component, but rather integrates the entire method without any modifications. The reason is that the occasional intra-class mixup does not conflict with our intra-class mixup method and provides modest additional benefits, such as increased diversity in training data and enhanced feature robustness.
@@
Our integrated solution has been extensively validated through experiments on six public image datasets, including CIFAR-100 [5], Food-101 [6], mini-Imagenet [7], OxfordIIIPet [8], and Caltech-256 [9]. MixUp and Manifold MixUp are selected as two implementations of the inter-class mixup component as they are seminal works in image-level and feature-level inter-class mixup, respectively. Experimental results demonstrate that the exclusive application of our intra-class mixup method yields an accuracy gain ranging from 0.22% to 3.09% (averaging at 1.25%) in 15 out of 18 experimental scenarios over the baseline devoid of any mixup operations. Furthermore, our integrated solution surpasses the best-performing method of either MixUp or our intra-class mixup method by an accuracy margin of 0.1% to 3.43%, with an average improvement of 1.16%. It also exceeds the top performer of either Manifold MixUp or our intra-class mixup method by 0.12% to 5.16%, with an average gain of 1.11%. These experimental results conclusively validate the effectiveness of our integrated solution in improving classification accuracy. Additionally, given that our integrated solution is model-agnostic, it holds a substantial opportunity for application in other domains like speech and text classification [1, 10] where MixUp or its variants have proven useful.
@@
    We have identified and underscored two critical limitations inherent in established mixup techniques such as MixUp and Manifold MixUp: a) neglect of intra-class mixup, leading to the underutilization of relationships among samples within the same class, and b) lack of improvement in intra-class cohesion through their mixing operations, limiting image classification performance.
@@
    We have proposed a significant enhancement to existing mixup methods with our innovative mixup technique. Our mixup technique effectively fortifies intra-class cohesion through intra-class mixing operations. This enhancement addresses the deficiencies found in many mixup approaches, providing an important complement to these methods. To the best of our knowledge, our method represents the first attempt into leveraging mixing operations to specifically improve intra-class cohesion.
@@
    We have developed an integrated solution specifically designed to augment both intra-class cohesion and inter-class separability through mixing operations. This approach effectively overcomes the two inherent limitations of existing mixup methods and represents a pioneering effort to harness mixing operations for these dual objectives.
@@
    We have introduced a new method for incorporating beneficial stochasticity into gradients. This method stands apart from existing methods such as random mini-batches, dropout, weight regularization, and data augmentation. Our findings suggest that the randomness of synthesized feature representations within local feature spaces contributes to the stochasticity of gradients and can enhance the accuracy of image classification. This discovery opens up a new avenue for research: exploring the impact of synthesized feature representation randomness on gradient stochasticity as a means for improving image classification performance.
The structure of this paper is as follows: Section 2 reviews related work. Section 3 details our proposed method. Section 4 presents the experimental results. Finally, Section 5 concludes this article.
@@ 2.2 Intra-class Mixup Technique
After reviewing the works mentioned above, it becomes evident that they inherently possess the following two significant limitations. The first is the neglect of intra-class mixup in contrast to inter-class mixup. The works presented in Section 2.1, as well as those applied to other domains beyond image classification (e.g., [10, 18, 19]), emphasize mixing between different classes while neglecting mixing within the same class. Although a few studies have specifically focused on intra-class mixup, they do not fully recognize its critical role in reducing intra-class variability. Furthermore, they do not leverage the combination of intra-class and inter-class mixup to maximize the model’s classification performance. The second limitation is the lack of augmenting the cohesion of image feature representations within the same class through their mixing operations. While these mixup methods succeed in improving inter-class separability via inter-class mixup, they do not address the enhancement of intra-class cohesion—a factor that is also crucial for image classification. To bridge these gaps, this article proposes a novel mixup method that strengthens the cohesion of intra-class feature representations, as well as a comprehensive integrated solution that improves both intra-class cohesion and inter-class separability.
@@ 3.1 Methodology
In the context of using data augmentation techniques with SGD, the loss calculation process of our integrated solution for a mini-batch is depicted in Figure 3. This process comprises four components:
@@
Figure 3: Pipeline of loss calculation for our integrated solution. It consists of four components including the supplementation component, intra-class mixup component, inter-class mixup component and integration component. $I_{1}$ through $I_{4}$ denote the four unaugmented original images, and $A_{1}$ through $A_{4}$ represent the augmented images obtained from $I_{1}$ through $I_{4}$ by applying a series of augmentation operations such as cropping, flipping, and cutout. Blue and orange colors are used to differentiate the two classes. Notably, many established mixup methods, such as MixUp and Manifold MixUp, can be used to implement the inter-class mixup component. Here, we only present the loss calculation process of MixUp for clarity.
@@ 3.1.1 Supplementation Component
Our intra-class mixup method, as the unique implementation for the intra-class mixup component, requires interpolating image feature representations from the same class within each mini-batch. This process requires that there be at least two images for each class contained in the mini-batch to facilitate the operation. SGD involves randomly sampling each mini-batch from the training dataset, which does not ensure that every class within the mini-batch will have a minimum of two images. Therefore, it becomes crucial to augment each mini-batch to guarantee that for every class appearing in the mini-batch, there are at least two images available.
@@ 3.1.2 Intra-class Mixup Component
Our intra-class mixup method is the exclusive implementation for the intra-class mixup component. Its details are as follows. For a given supplemented mini-batch sequence ($I_{1}$, $y_{1}$), …, ($I_{\hat{N}}$, $y_{\hat{N}}$) where $y_{i}$ represents the class label of the unaugmented original image $I_{i}$ and $\hat{N}$ the supplemented mini-batch size, we input $I_{1},...,I_{\hat{N}}$ into the model and obtain the vectors $\bm{f}_{1},...,\bm{f}_{\hat{N}}$ fed into the final classification layer. These vectors serve as the feature representations of these images. The single synthesized feature representation $\bm{f}_{\text{syn}}^{c}$ for the class labeled $c$ is defined as:
@@ 3.2 Method Explanation
This subsection thoroughly explains the rationale behind our integrated solution by addressing the following three pivotal questions. Additionally, we detail the development of the loss calculation formula, tracing its evolution from the initial formulation in Eq. (6) to its finalized version in Eq. (5). To provide a clear response to the first question, we assume that the feature representations of augmented images, rather than those of unaugmented original images, are employed to generate synthetic feature representations. In relation to the second question, we will explore the potential side effect of this assumption and examine why utilizing the feature representations of unaugmented original images to produce synthetic representations can alleviate this side effect.
(1) Why can our intra-class mixup method improve intra-class cohesion? As shown in Figure (a), each randomly synthesized feature representation can be viewed as a point within the convex hull or on a line segment formed by the feature representations of augmented images that generate it. Moreover, the last fully-connected layer, which receives image feature representations as input, acts as a linear classifier. In this case, given the requirement that any point within the geometric space formed by feature representations of augmented images from the same class must be correctly classified, these feature representations are encouraged to cluster more tightly together, thereby expediting the fulfillment of this requirement. Crucially, the presence of this encouragement is consistent throughout the training process. As a result, compared to the scenario where this encouragement is absent, there is not only an enhancement in intra-class cohesion upon completion of training, but also a more rapid aggregation of same-class image feature representations during the training process, leading to consistently higher classification accuracy, as demonstrated in Section 4.2.1. In the implementation of intra-class mixup, we ensure that each weight $w_{j}\geq 0$, which guarantees that the synthesized feature representations lie within the formed geometric space. By generating $w_{j}$ via random numbers sampled from a uniform distribution $U(0,1)$, we ensure the complete randomness of synthesized representations within the formed geometric space.
@@
(2) Why is it beneficial to leverage feature representations from unaugmented original images in order to generate synthesized ones? The reasonableness of expecting synthesized feature representations to be correctly classified hinges on the assumption that they accurately reflect the true distribution. However, this assumption does not always hold. Figure (b) clearly demonstrates that even when a synthesized representation is generated from authentic feature representations, it does not belong to the true distribution. Furthermore, the negative impact will be amplified when it overlaps with the feature distribution of another class.
@@
(3) What are the benefits of combining intra-class mixup with inter-class mixup? In our integrated solution, the intra-class mixup method works synergistically with the inter-class mixup method. The former ensures that image feature representations of the same class are more tightly clustered. Meanwhile, the latter method increases inter-class separability, making different classes more distinguishable from one another. The seamless integration of these methods yields a comprehensive loss function that incorporates both intra- and inter-class mixup strategies in a harmonious balance. Consequently, this dual integration serves to concurrently improve intra-class cohesion and inter-class separability. Moreover, the introduction of the hyperparameter $\beta$ allows for an adjustable balance, which is essential for training a model that is well-suited for the complexities of image classification. This adaptability enables the model to account for the unique commonalities within classes and distinctions among classes inherent in different image classification tasks. Building on these benefits, Eq. (8) is augmented by substituting its $\mathscr{L}_{\text{augment}}$ with $\mathscr{L}_{\text{inter}}$, culminating in the final loss equation presented in Eq. (5).

### 24 Mar 2024 v2

# SynerMix: Synergistic Mixup Solution for Enhanced Intra-Class Cohesion and Inter-Class Separability in Image Classification

@@ Abstract
To address the issues of MixUp and its variants (e.g., Manifold MixUp) in image classification tasks—namely, their neglect of mixing within the same class (intra-class mixup) and their inadequacy in enhancing intra-class cohesion through their mixing operations—we propose a novel mixup method named SynerMix-Intra and, building upon this, introduce a synergistic mixup solution named SynerMix. SynerMix-Intra specifically targets intra-class mixup to bolster intra-class cohesion, a feature not addressed by current mixup methods. For each mini-batch, it leverages feature representations of unaugmented original images from each class to generate a synthesized feature representation through random linear interpolation. All synthesized representations are then fed into the classification and loss layers to calculate an average classification loss that significantly enhances intra-class cohesion. Furthermore, SynerMix combines SynerMix-Intra with an existing mixup approach (e.g., MixUp, Manifold MixUp), which primarily focuses on inter-class mixup and has the benefit of enhancing inter-class separability. In doing so, it integrates both inter- and intra-class mixup in a balanced way while concurrently improving intra-class cohesion and inter-class separability. Experimental results on six datasets show that SynerMix achieves a 0.1% to 3.43% higher accuracy than the best of either MixUp or SynerMix-Intra alone, averaging a 1.16% gain. It also surpasses the top-performer of either Manifold MixUp or SynerMix-Intra by 0.12% to 5.16%, with an average gain of 1.11%. Given that SynerMix is model-agnostic, it holds significant potential for application in other domains where mixup methods have shown promise, such as speech and text classification. Our code is publicly available at: https://github.com/wxitxy/synermix.git.
@@ 1 Introduction
Figure 2: t-SNE visualization of feature representations for testing images from the CIFAR-100 dataset: comparing three methods – the baseline (no mixing operation), Manifold MixUp, and SynerMix-Intra.
@@
Despite the significant success achieved by existing mixup techniques, a closer examination reveals two critical limitations. The first concerns the mixing strategies—namely, intra-class and inter-class mixup. Existing mixup methods such as MixUp, Manifold MixUp, CutMix, and SaliencyMix often neglect intra-class mixup, which involves blending images or hidden representations from the same class. This oversight leads to an underutilization of the relationships among samples within the same class, potentially constraining image classification performance. The root of this issue lies in the shuffle operation; it results in a disproportionately high number of mixed pairs generated from different classes compared to those from the same class. Consequently, there is a predominant emphasis on inter-class mixup, which mixing images or hidden representations from different classes. As illustrated in Figure 1, with 16 classes and a mini-batch size of 32, the average probability of occurrence for intra-class mixup during a shuffle operation is only approximately 9%. If the number of classes and the mini-batch size both increase to 128, this probability decreases to below 2%. In this article, we refer to these techniques that primarily focus on inter-class mixup as “inter-class mixup methods.”
@@
Additionally, the second limitation arises with respect to the mixing outcomes. Existing mixup methods, such as MixUp and Manifold MixUp, fall short in enhancing intra-class cohesion of image feature representations through mixing operations. In this article, the feature representation of an image refers to the input vector that is fed into the neural network’s final classification layer—a fully connected layer responsible for predicting scores for each class. To be specific, the incidental intra-class mixup in current mixup methods indeed has the potential to facilitate the model’s learning of common features among images from the same class, potentially leading to more cohesive clustering of feature representations. Nevertheless, the goal of forming compact class clusters is not explicit, and its effectiveness is significantly diminished due to the rarity of intra-class mixup. On the other hand, the predominant inter-class mixup they implement does enhance the distinction among different classes by promoting smoother predictions between classes but fails to offer any mechanism for reducing intra-class variability. As illustrated in Figure 2, especially within the region enclosed by the dashed blue box, the distinction among classes achieved by Manifold MixUp is noticeably improved compared to the baseline devoid of any mixing operations. However, the feature distribution of each class does not show a marked increase in clustering compared to the baseline. This is particularly obvious in comparison to the enhanced cohesion achieved by our mixup method named SynerMix-Intra proposed in this article, as is evident within the red dashed box. In classification tasks, both intra-class cohesion and inter-class separability are known to be vital and equally significant. Consequently, the lack of a dedicated mechanism to improve intra-class cohesion in existing mixup methods may limit their performance in classification tasks.
@@
To overcome these limitations, we propose the novel mixup method SynerMix-Intra that significantly enhances intra-class cohesion through targeted intra-class mixup. Furthermore, we introduce a synergistic mixup solution named SynerMix that effectively combines the benefits of both SynerMix-Intra and existing mixup methods, addressing the two critical limitations identified in current mixup techniques.
@@
SynerMix-Intra focuses on intra-class mixup, an aspect that existing mixup techniques overlook, with the aim of enhancing intra-class cohesion—a benefit not provided by current mixup techniques. Specifically, during training, it generates a single synthesized feature representation for each class in every mini-batch. This synthesis is achieved through random linear interpolation among the feature representations of unaugmented original images within the same class. After the synthesized feature representations for all classes in the mini-batch are generated, they are fed into the final classification layer to produce predictions. An intra-class mixup loss is then calculated by averaging the classification losses of these synthesized representations. This loss is designed to promote the model to cluster image feature representations from the same class more tightly, resulting in enhanced intra-class cohesion at the end of training.
@@
Building upon SynerMix-Intra and existing mixup approaches, we have developed the synergistic mixup solution SynerMix. SynerMix leverages an existing mixup method—such as MixUp or Manifold MixUp—as the inter-class mixup component, while SynerMix-Intra serves as the intra-class counterpart. The intra- and inter-class mixup components work together in a complementary manner, with the final loss calculated as a weighted sum of their losses, modulated by a balancing hyperparameter $\beta$. In doing so, SynerMix not only uniquely combines both inter- and intra-class mixup strategies in a balanced manner, but also simultaneously improves both intra-class cohesion and inter-class separability. It is capable of adapting to the distinct commonalities within classes and differences among classes that are inherent in various image classification tasks, thereby enhancing the model’s generalization capabilities and robustness. Significantly, SynerMix allows for flexible substitution of the inter-class mixup component; numerous existing mixup methods, including image-level mixup techniques like MixUp and feature-level mixup techniques like Manifold MixUp, can be implemented as the inter-class mixup component.
@@
It is worth noting that SynerMix does not deliberately exclude the occasional intra-class mixup that occurs in the existing mixup method, but rather integrates the entire method without any modifications. The reason is that the occasional intra-class mixup does not conflict with SynerMix-Intra and provides modest additional benefits, such as increased diversity in training data and enhanced feature robustness.
@@
SynerMix has been extensively validated through experiments on six public image datasets, including CIFAR-100 [5], Food-101 [6], mini-Imagenet [7], OxfordIIIPet [8], and Caltech-256 [9]. MixUp and Manifold MixUp are selected as two implementations of the inter-class mixup component as they are seminal works in image-level and feature-level inter-class mixup, respectively. Experimental results demonstrate that the exclusive application of SynerMix-Intra yields an accuracy gain ranging from 0.22% to 3.09% (averaging at 1.25%) in 15 out of 18 experimental scenarios over the baseline devoid of any mixup operations. Furthermore, SynerMix surpasses the best-performing method of either MixUp or SynerMix-Intra by an accuracy margin of 0.1% to 3.43%, with an average improvement of 1.16%. It also exceeds the top performer of either Manifold MixUp or SynerMix-Intra by 0.12% to 5.16%, with an average gain of 1.11%. These experimental results validate the effectiveness of SynerMix in concurrently improving intra-class cohesion and inter-class separability. Additionally, given that SynerMix is model-agnostic, it holds a substantial opportunity for application in other domains like speech and text classification [1, 10] where MixUp or its variants have proven useful.
@@
    We have identified and underscored two critical limitations inherent in established mixup approaches such as MixUp and Manifold MixUp: a) neglect of intra-class mixup, leading to the underutilization of relationships among samples within the same class, and b) inadequacy in improving intra-class cohesion through their mixing operations, limiting image classification performance.
@@
    We have proposed a significant enhancement to existing mixup techniques with our mixup method SynerMix-Intra. SynerMix-Intra effectively fortifies intra-class cohesion through intra-class mixing operations. This enhancement addresses the deficiencies found in numerous mixup approaches, providing an important complement to these methods. To the best of our knowledge, SynerMix-Intra represents the first attempt into leveraging mixing operations to specifically improve intra-class cohesion.
@@
    We have developed the synergistic mixup solution SynerMix specifically designed to augment both intra-class cohesion and inter-class separability through mixing operations. This approach effectively overcomes the two inherent limitations of existing mixup methods and represents a pioneering effort to harness mixing operations for these dual objectives.
@@
    We have found a new method for incorporating beneficial stochasticity into gradients. This method stands apart from existing methods such as random mini-batches, dropout, weight regularization, and data augmentation. Our findings suggest that the randomness of synthesized feature representations within local feature spaces contributes to gradient stochasticity and can enhance the accuracy of image classification. This discovery opens up a new avenue for research: exploring the impact of synthesized feature representation randomness on gradient stochasticity as a means for improving image classification performance.
The structure of this paper is as follows: Section 2 reviews related work. Section 3 details the synergistic mixup solution SynerMix. Section 4 presents the experimental results. Finally, Section 5 concludes this article.
@@ 2.2 Intra-class Mixup Technique
After reviewing the works mentioned above, it becomes evident that they inherently possess the following two significant limitations. The first is the neglect of intra-class mixup in contrast to inter-class mixup. The works presented in Section 2.1, as well as those applied to other domains beyond image classification (e.g., [10, 18, 19]), emphasize mixing between different classes while neglecting mixing within the same class. Although a few studies have specifically focused on intra-class mixup, they do not fully recognize its critical role in reducing intra-class variability. Furthermore, they do not leverage the combination of intra-class and inter-class mixup to maximize the model’s classification performance. The second limitation is the lack of augmenting the cohesion of image feature representations within the same class through their mixing operations. While these mixup methods succeed in improving inter-class separability via inter-class mixup, they do not address the enhancement of intra-class cohesion—a factor that is also crucial for image classification. To bridge these gaps, this article proposes the novel mixup method SynerMix-Intra that strengthens the cohesion of intra-class feature representations, as well as the synergistic mixup solution SynerMix that improves both intra-class cohesion and inter-class separability.
@@ 3.1 Methodology
In the context of using data augmentation techniques with SGD, the loss calculation process of SynerMix for a mini-batch is depicted in Figure 3. This process comprises four components:
@@
Figure 3: Pipeline of loss calculation for SynerMix. It consists of four components including the supplementation component, intra-class mixup component, inter-class mixup component and integration component. $I_{1}$ through $I_{4}$ denote the four unaugmented original images, and $A_{1}$ through $A_{4}$ represent the augmented images obtained from $I_{1}$ through $I_{4}$ by applying a series of augmentation operations such as cropping, flipping, and cutout. Blue and orange colors are used to differentiate the two classes. Notably, many established mixup methods, such as MixUp and Manifold MixUp, can be used to implement the inter-class mixup component. Here, we only present the loss calculation process of MixUp for clarity.
@@ 3.1.1 Supplementation Component
SynerMix-Intra, as the unique implementation for the intra-class mixup component, requires interpolating image feature representations from the same class within each mini-batch. This process requires that there be at least two images for each class contained in the mini-batch to facilitate the operation. SGD involves randomly sampling each mini-batch from the training dataset, which does not ensure that every class within the mini-batch will have a minimum of two images. Therefore, it becomes crucial to augment each mini-batch to guarantee that for every class appearing in the mini-batch, there are at least two images available.
@@ 3.1.2 Intra-class Mixup Component
SynerMix-Intra is the exclusive implementation for the intra-class mixup component. Its details are as follows. For a given supplemented mini-batch sequence ($I_{1}$, $y_{1}$), …, ($I_{\hat{N}}$, $y_{\hat{N}}$) where $y_{i}$ represents the class label of the unaugmented original image $I_{i}$ and $\hat{N}$ the supplemented mini-batch size, we input $I_{1},...,I_{\hat{N}}$ into the model and obtain the vectors $\bm{f}_{1},...,\bm{f}_{\hat{N}}$ fed into the final classification layer. These vectors serve as the feature representations of these images. The single synthesized feature representation $\bm{f}_{\text{syn}}^{c}$ for the class labeled $c$ is defined as:
@@ 3.2 Method Explanation
This subsection thoroughly explains the rationale behind SynerMix by addressing the following three pivotal questions. Additionally, we detail the development of the loss calculation formula, tracing its evolution from the initial formulation in Eq. (6) to its finalized version in Eq. (5). To provide a clear response to the first question, we assume that the feature representations of augmented images, rather than those of unaugmented original images, are employed to generate synthetic feature representations. In relation to the second question, we will explore the potential side effect of this assumption and examine why utilizing the feature representations of unaugmented original images to produce synthetic representations can alleviate this side effect.
(1) Why can SynerMix-Intra improve intra-class cohesion? As shown in Figure 4(a), each randomly synthesized feature representation can be viewed as a point within the convex hull or on a line segment formed by the feature representations of augmented images that generate it. Moreover, the last fully-connected layer, which receives image feature representations as input, acts as a linear classifier. In this case, given the requirement that any point within the geometric space formed by feature representations of augmented images from the same class must be correctly classified, these feature representations are encouraged to cluster more tightly together, thereby expediting the fulfillment of this requirement. Crucially, the presence of this encouragement is consistent throughout the training process. As a result, compared to the scenario where this encouragement is absent, there is not only an enhancement in intra-class cohesion upon completion of training, but also a more rapid aggregation of same-class image feature representations during the training process, leading to consistently higher classification accuracy, as demonstrated in Section 4.2.1. In the implementation of intra-class mixup, we ensure that each weight $w_{j}\geq 0$, which guarantees that the synthesized feature representations lie within the formed geometric space. By generating $w_{j}$ via random numbers sampled from a uniform distribution $U(0,1)$, we ensure the complete randomness of synthesized representations within the formed geometric space.
@@
(2) Why is it beneficial to use feature representations from unaugmented original images in order to generate synthesized ones? The reasonableness of expecting synthesized feature representations to be correctly classified hinges on the assumption that they accurately reflect the true distribution. However, this assumption does not always hold. Figure 4(b) clearly demonstrates that even when a synthesized representation is generated from authentic feature representations, it does not belong to the true distribution. Furthermore, the negative impact will be amplified when it overlaps with the feature distribution of another class.
@@
(3) What are the benefits of combining intra-class mixup with inter-class mixup? In SynerMix, the intra-class mixup component works synergistically with the inter-class mixup component. The former ensures that image feature representations of the same class are more tightly clustered. Meanwhile, the latter increases inter-class separability, making different classes more distinguishable from one another. The seamless integration of these methods yields a comprehensive loss function that incorporates both intra- and inter-class mixup strategies in a harmonious balance. Consequently, this dual integration serves to concurrently improve intra-class cohesion and inter-class separability. Moreover, the introduction of the hyperparameter $\beta$ allows for an adjustable balance, which is essential for training a model that is well-suited for the complexities of image classification. This adaptability enables the model to account for the unique commonalities within classes and distinctions among classes inherent in different image classification tasks. Building on these benefits, Eq. (8) is augmented by substituting its $\mathscr{L}_{\text{augment}}$ with $\mathscr{L}_{\text{inter}}$, culminating in the final loss equation presented in Eq. (5).